# FedSEA-LLaMA: A Secure, Efficient and Adaptive Federated Splitting Framework for Large Language Models

### v3

@@ Introduction
While the success of Large Language Models (LLMs) has largely stemmed from their ability to process vast amounts of public data (Kaplan et al. 2020), the true potential of these models lies in their application to high-quality private data. Private datasets often contain more granular and specific insights about individuals or organizations, typically collected through more controlled and precise methodologies (Yang et al. 2024). Additionally, private data tend to be more timely, targeted, and unique, thus providing models with semantic contexts and environmental settings that more closely mirror real-world scenarios (Ye et al. 2024). Therefore, effectively incorporating private data will further advance the performance of LLMs.
@@
The distributed private data silos and the significant computational demands of LLMs pose substantial challenges for their deployment in federated environments. Private data is often isolated in local silos, such as mobile devices (Ghimire and Rawat 2022), enterprise servers (Huang et al. 2024), or medical institutions (Yuan et al. 2024), as centralized storage and processing of such data may raise privacy concerns and regulatory risks (Chen et al. 2024). Federated Learning (FL) (Stremmel and Singh 2021) offers a promising solution by enabling local training on user devices and only sending model parameters for aggregation. They often leverage Adapters (Houlsby et al. 2019) or LoRA (Hu et al. 2022) as transmitted parameters, or employ collaborative training between large and small models (Hinton, Vinyals, and Dean 2015) to fine-tune LLMs. However, in extremely resource-constrained scenarios (e.g., less than 3.5GB of GPU memory), fine-tuning the LLMs remains infeasible. Therefore, how to effectively deploy LLMs in FL frameworks while ensuring privacy and scalability in resource-constrained scenarios remains an open research problem.
To address this, transformer-based federated splitting models (Zheng et al. 2024; Su et al. 2024) have been proposed to adapt LLM fine-tuning to resource-constrained scenarios. Titanic (Su et al. 2024) handles limited server resources by partitioning LLM across clients and enabling collaborative training. FL-GLM (Zheng et al. 2024) offloads most model parameters to a resource-ample server while retaining the input and output layer on the client. As shown in Figure 1a, FL-GLM divides Transformers into three parts, i.e., the first input block and the last output block on the client, while the middle blocks on the server. The forward propagation (FP) path is client-to-server-to-client with peer-to-peer key encryption. After the client calculates loss, gradients backpropagate (BP) from client-to-server-to-client with the same encryption. Although FL-GLM provides a potentially viable solution for clients in resource-constrained scenarios, it still faces several limitations: 1) Privacy Leakage: during forward and backward propagation, activation values and gradients from the input blocks are transmitted, which can be exploited to reconstruct original input data through reverse engineering (Asnani et al. 2023). 2) Communication Cost: LLMs are auto-regressive, requiring sequential token generation based on prior tokens. In federated split learning, this necessitates a full forward and backward pass with client-server communication for each token, greatly increasing communication rounds and bandwidth usage. 3) Lack of Adaptability: Fixed partition points of FL-GLM lack flexibility to downstream tasks and hardware constraints, as different devices vary in their capacity to handle workloads and parameter sizes.
@@
Traditional Federated Learning (FL) frameworks face scalability challenges with large language models (LLMs) due to high communication and computational costs (McMahan et al. 2017; Stremmel and Singh 2021; Ji et al. 2019). Recent federated LLM frameworks (Fan et al. 2023; Kuang et al. 2024; Ye et al. 2024) address this by using instruction tuning, Adapters, and LoRA to reduce parameter updates, lowering communication overhead. However, transmission demands often still exceed system limits. To mitigate this, gradient compression (Wu et al. 2024b; Shu et al. 2024) and quantization (JianHao et al. 2024) techniques have been applied. Structural pruning based on Fisher Information (Liu et al. 2024) further narrows updates to key layers, reducing model size without performance loss. Knowledge distillation (Wu et al. 2024a; Fan et al. 2024, 2025) offers another solution, using smaller proxy models to transfer knowledge with minimal communication. Client heterogeneity remains a major challenge (Zhang et al. 2024), prompting methods like LoRA-based stacking (Bai et al. 2024; Wang et al. 2024) to align models from diverse clients and improve fairness.
Despite advances in efficient fine-tuning and parameter reduction for federated LLMs, resource-constrained clients still face challenges with independent fine-tuning, particularly under compression overhead. Split learning (Thapa et al. 2022) offers a promising alternative in such settings.
@@
Split learning is an emerging decentralized training approach designed for resource-constrained settings (Abedi and Khan 2024; Rahman et al. 2020; Matsubara and Levorato 2021), especially mobile or GPU-limited clients. It works by dividing a neural network into sub-networks, each processed on a different device for collaborative learning. To address the training of sequential data in LLMs, FedBERT (Tian et al. 2022) introduces an innovative federated learning framework designed to pre-train language models in a split architecture, addressing the challenge of limited computational capacity on client devices. Although this design enables more efficient distribution of training workloads, it comes with drawbacks, such as increased communication overhead, fixed partition points, and susceptibility to privacy threats like embedding gradient attacks.
Recently, Su et al. (2024) extends the scope of split learning by exploring scenarios with heterogeneous resources, particularly considering a case where the server lacks sufficient computational capacity. In their approach, the LLM is automatically divided into $k$ partitions, each assigned to one of $k$ clients, while the server merely performs aggregation. Although this strategy alleviates server-side computation, it considerably increases the overall training and inference time. Furthermore, repeated transmission of hidden states increases the risk of single points of failure and introduces potential security vulnerabilities. Zheng et al. (2024) propose FL-GLM to place the input and output blocks locally on client devices, while the remaining primary model parameters are centralized on a server with ample computational resources. And then, they employ key encryption during client-server communication to prevent reverse engineering attacks from peer clients. But the peer-to-peer key encryption is not enough because the peer clients and servers can collaborate to infer the clients’ private data.
@@
We design an extreme attack scenario to evaluate the robustness of FedSEA-LLaMA. Inspired by Pasquini, Ateniese, and Bernaschi (2021), we construct a model inversion attack in a multi-client federated setting, as illustrated in Figure 3a, where a malicious client colludes with an honest-but-curious server. The malicious client shares its private data in plaintext with the server, allowing the server to train an attack model locally. The goal of the attack model is to reconstruct the original input text from the hidden states transmitted by the client. During this process, the server does not tamper with the hidden states sent to other honest clients, thereby performing a stealthy inversion attack against their private data without awareness.
@@
During backpropagation, according to (Clark 2017), the gradient of $W_{n}$ is:
@@
Model ReCoRD COPA WSC RTE BoolQ WiC CB MultiRC CoQA Xsum F1/Acc. Acc. Acc. Acc. Acc. Acc. F1 F1a/EM F1/EM rouge-1/rouge-2 $\text{T5}_{\text{large}}$ (Raffel et al. 2020) 85.7/85.0 78.0 84.6 84.8 84.3 71.6 96.4 80.9/46.6 - 40.9/17.3 $\text{BART}_{\text{Large}}$ (Lewis et al. 2020) 88.3/87.8 60.0 65.4 84.5 84.3 69.0 90.5 81.8/48.0 - 45.1/22.3 $\text{GLM}_{\text{RoBERTa}}$ (Du et al. 2022) 89.6/89.0 82.0 83.7 87.7 84.7 71.2 98.7 82.4/50.1 - 45.5/23.5 ChatGLM-6B (Zeng et al. 2022) 80.2/78.7 85.0 71.2 81.6 83.4 71.0 85.7 78.2/45.6 - 37.6/12.5 FL-GLM (Zheng et al. 2024) 79.8/78.4 85.0 71.2 80.1 81.9 69.6 85.7 79.3/46.1 62.7/49.0 37.0/11.9 LLaMA2-7B (Touvron et al. 2023) 81.3/79.5 75.0 68.3 73.3 83.8 70.9 85.7 82.3/50.1 88.6/80.9 45.9/26.1 FedSEA-LLaMA 81.3/79.4 75.0 68.4 73.3 82.4 70.8 85.7 82.7/48.0 88.6/80.7 47.6/25.0
@@
  • The SuperGLUE benchmark (Wang et al. 2020) evaluates eight distinct NLU tasks, from common sense reasoning to fine-grained semantic interpretation.
  • The CoQA dataset (Reddy, Chen, and Manning 2019) is designed for the task of conversational question answering(QA), featuring 8k+ dialogues across seven domains. Answers are free-form text, with nearly half of the questions requiring coreference resolution and pragmatic reasoning, simulating real-world dialogue coherence.
  • The XSum dataset (Narayan, Cohen, and Lapata 2018) evaluates abstractive summarization via extreme compression of BBC news articles into single-sentence summaries. It tests the model’s ability to abstract and condense salient information across diverse topics.
@@
For conversational QA tasks, we use Exact Match (EM) and F1 score to evaluate model performance (Yatskar 2019; You et al. 2022). EM measures the percentage of predictions that exactly match the ground truth, accounting for punctuation and case, while F1 reflects the harmonic mean of precision and recall, rewarding partial overlaps.
For summarization tasks, we use standard ROUGE metrics (Lin 2004; Liu et al. 2021; Fang et al. 2022), specifically ROUGE-1 and ROUGE-2, which assess unigram and bigram overlap between generated and reference summaries, capturing content coverage and fluency.
@@
We evaluate FedSEA-LLaMA based on the LLaMA2-7B model (Touvron et al. 2023), an open-source pre-trained model developed by Meta and FL-GLM. We additionally include several strong baselines: $\text{T5}_{\text{large}}$ (Raffel et al. 2020), $\text{BART}_{\text{Large}}$ (Lewis et al. 2020), $\text{GLM}_{\text{RoBERTa}}$ (Du et al. 2022) and ChatGLM-6B (GLM et al. 2024).
@@
The server is equipped with a Gigabit Ethernet card and utilizes multiple GPUs, including two NVIDIA A6000 and two NVIDIA L40. The implementation is based on the Flower (Beutel et al. 2020) framework, which orchestrates the transfer of intermediate data between clients and the server. Due to limited computational resources, FP32 precision and a batch size of 1 are used only for the CoQA dataset. All other experiments are conducted with FP16 precision and a batch size of 2. The random seed is consistently set to 42. Fine-tuning is performed using distributed LoRA training, where $Blocks_{A}$, $Blocks_{B}$, and $Blocks_{C}$ each maintain their own LoRA adapters. For all LoRA fine-tuning experiments in this paper, the target modules are set to q_proj, v_proj, k_proj, and o_proj.
@@ Acknowledgements
Appendix A Noise Added to Hidden States
(a) Loss Curves for Noise Scales
($Blocks_{A}$=1, $Blocks_{C}$=1)
(b) Loss Curves for Noise Scales
($Blocks_{A}$=2, $Blocks_{C}$=1)
(c) Loss Curves for Noise Scales
($Blocks_{A}$=3, $Blocks_{C}$=1)
Figure 6: FedSEA-LLaMA: Noise Injection across different number of blocks in $Blocks_{A}$ on MultiRC dataset.
The magnitude of noise injection significantly influences the stability and convergence speed of model training. Given that noise is directly injected into the hidden state $h_{A}$, the relative magnitude between the noise and $h_{A}$ naturally affects the model’s performance. To quantify this effect, we first fine-tune the FedSEA-LLaMA model on the MultiRC (Khashabi et al. 2018) with a batch size of 2 and a learning rate of $1\times 10^{-5}$, analyzing the average absolute magnitude of all elements within the hidden state $h_{A}$. According to Figure 7, the average magnitude of the hidden state $h_{A}$ is approximately 0.02 ($num_{Blocks_{A}}$=1). The average magnitude of the hidden state $h_{A}$ increases with the number of blocks in $Blocks_{A}$.
Based on this observation, we subsequently inject noise at various scales around this value during training to investigate how the noise magnitude affects both the convergence speed and downstream task performance. Training is conducted for 5,000 steps with FP16 precision. We use LoRA with a rank of 8 and a scaling factor $\alpha=16$. The learning rate is 1e-5, and the batch size is 2. Corresponding experimental results are presented in Figure 6(a)-6(c) and Table 2. To assess the sensitivity to the noise scale when $Blocks_{A}$=3, we evaluate performance across a range of noise scales (from 0.1 to 0.4). As shown in Table 5, model performance decreases as the noise scale increases. This suggests that while moderate noise injection is tolerable, excessive noise harms representation quality.
Noise Scale F1a EM
0.1 75.68 34.10
0.15 71.21 26.82
0.2 59.08 16.47
0.3 55.79 7.45
0.4 57.68 6.09
Table 5: Performance under different noise scales when $Blocks_{A}$=3
A horizontal comparison reveals that, under the same noise level, increasing the number of blocks in $Blocks_{A}$ consistently accelerates convergence. Table 2 indicates that when the noise amplitude reaches 0.1, only the single-block configuration of $Blocks_{A}$ shows a significant drop in downstream performance, whereas the three-block version of $Blocks_{A}$ demonstrates greater robustness. This can be explained by the average magnitude shown in Figure 7: as the depth of $Blocks_{A}$ increases, the ratio between the average absolute value of the hidden state $h_{A}$ and the noise amplitude also increases, thereby improving the signal-to-noise ratio (SNR). This enhanced SNR is precisely what contributes to the better convergence and stronger robustness observed with a larger number of $Blocks_{A}$.
Figure 7: Avg magnitude of hidden state $h_{A}$ ($Blocks_{C}$ =1) on MultiRC dataset.
Appendix B Prompt
Table 6 illustrates examples of cloze-style questions and corresponding answers derived from each task. (1)ReCoRD tests reading comprehension with common sense reasoning by resolving ambiguous entities in passages. (2)COPA assesses causal reasoning by choosing the more likely cause or effect for a given premise. (3)WSC focuses on pronoun disambiguation requiring deep contextual understanding. (4)RTE evaluates textual entailment by determining if one sentence follows, contradicts, or is unrelated to another. (5)BoolQ involves yes/no questions answered using evidence from context passages. (6)WiC tests word sense disambiguation by checking if a word has the same meaning in two sentences. (7)CB deals with entailment in complex sentences with nuanced reasoning. (8)MultiRC features multi-answer questions based on multi-sentence contexts, requiring integration of dispersed information.
Table 7 illustrates examples of fine-tuning templates of CoQA and XSum.
Dataset Task Cloze Question Answers ReCoRD Question answering [passage p] [cloze question q] Answer candidates COPA Causal reasoning “[choice c1]” or“[choice c2]”? [premise p], so [M]. c1/c2 WSC Coreference resolution [sentence s] The pronoun ‘*p*’ refers to [M]. Noun n RTE Textual entailment “[hypothesis h]”? [M] “[premise p]” “yes”/“no” BoolQ Question answering [passage p]. Question: q? Answer:[M]. “yes” / “no” WiC Word sense disambiguation “[sentence s1]”/“[sentence s2]”Similar sense of [word w]? [M]. “yes”/“no” CB Textual entailment “[hypothesis h]”? [M], “[premise p]” “yes”/“no”/“maybe” MultiRC Question answering [passage p]. Question: q? Is it [answer a]? [M]. “yes”/“no”
Table 6: Cloze questions and answers for the 8 SuperGLUE tasks
Dataset Task Fine-tuning template
CoQA conversational QA
<s><s>[INST] <<SYS>><<SYS>>You are a helpful assistant.
You need to answer my question base on the provided story.<</SYS>><</SYS>>
The story is {story}.My question is {question}.[/INST]{answer}</s></s>
XSum summarization generation
<s><s>[INST] <<SYS>><<SYS>>You are a helpful assistant.
You need to summarize an abstract base on the provided passage.<</SYS>><</SYS>>
The passage is {passage}.[/INST]{summarization} </s></s>
Table 7: Fine-tuning templates for CoQA and XSum
Appendix C Parallel Training
[…]
Table 8: Comparison of training time between different training strategies on ReCoRD dataset.
(a) Client-batch parallel training.
(b) Server-hierachical parallel training.
Figure 8: Multi-client Parallel Training of FedSEA-LLaMA. In (a), the server receives forward-pass states from multiple clients simultaneously, concatenates them into a batch, and performs joint forward propagation. Then, the server splits hidden state $h_{B}$ and sends them back to different clients. In (b), each $client_{i}$ is paired with a corresponding $sub\_server_{i}$. After several training steps, the central server aggregates all sub-servers’ models using FedAvg and distributes the updated global server model back to each $sub\_server_{i}$.
In multi-client settings, the FedSEA-LLaMA framework accommodates three training paradigms: sequential training, client-batch parallelism, and server-hierarchical parallelism, each designed for different system configurations and resource availabilities.
Sequential Training
The client begins by passing hidden states through its input blocks $Blocks_{A}$ and forwards the result to the server. The server processes the input through $Blocks_{B}$ and returns the output states. The client then computes the loss with $Blocks_{C}$, sends the gradients to the server for model updates, and receives the updated gradients to adjust its input blocks. This process continues until all clients have completed their respective training phases for the current round. Although this method is inherently slower due to its serialized nature, it significantly reduces the computational burden on the server, making it a suitable choice for scenarios where server-side resources are limited.
Client-batch Parallelism
The client-batch parallel approach enhances training throughput by expanding the effective batch size at the server, as shown in Figure  8(a). During each training round, the server collects intermediate hidden states from all participating clients. If each of the $M$ clients processes a local batch of size $b$, the hidden states sent to the server will each have the shape $(b,\textit{seq\_length},\textit{hidden\_size})$. The server concatenates these hidden states along the batch dimension, forming a combined tensor of shape $(b\times M,\textit{seq\_length},\textit{hidden\_size})$. This aggregated batch is then used to perform a joint forward and backward pass on the server model, allowing it to compute gradients more efficiently and update shared parameters.
Server-hierarchical Parallelism
In a server-hierarchical parallel setup, the central server spawns multiple sub-server model instances, each dedicated to a specific client, illustrated in Figure  8(b). These instances run in parallel, enabling simultaneous local training for all clients. Typically, after a fixed number of local updates, the central server synchronizes the training progress by aggregating model parameters across all sub-servers. This is usually done via weighted averaging of both server-side and client-side model parameters. This strategy enables significant improvements in training efficiency and scalability, especially in environments where the server has access to multiple GPUs or multi-threaded execution capabilities.
For communication cost analysis, we randomly select 1,000 samples from the ReCoRD dataset. All training uses a batch size of 1 with FP16 precision for efficiency. LoRA is applied with a rank of 8 and a scaling factor of 16, targeting attention modules during fine-tuning. The model is partitioned with $Blocks_{A}$=1 and $Blocks_{C}$=1. Each experiment is repeated five times, and we report the mean and standard deviation of the total communication time in Table 8. Results show that sequential training is the slowest, client-batch training is faster, and server hierarchical training is the fastest. Notably, when the server can handle larger batches, increasing the number of clients in client-batch training effectively reduces training time.
Appendix D Efficiency Comparison between FedSEA-LLaMA and FL-GLM
For the measurements of inference speed, inference memory usage, and per-token communication volume, we used the following prompt template:
{B_INST}{B_SYS}System: You are a helpful assistant. Always answer as helpfully as possible.{E_SYS}User: what can you do?{E_INST}
The total prompt length was 44 tokens. All inference experiments were conducted in FP16 precision with a maximum output length of 128 tokens. During inference, the client held one layer of model A and one layer of model C.
For the training memory usage experiments, we used the MultiRC dataset. Both model A and model C contained one layer. Unless otherwise specified, the default configuration was as follows: learning rate $1\times 10^{-5}$, LoRA rank =32=32, LoRA scaling factor =32=32, random seed =64=64, FP16 precision, and batch size =2=2. It is important to note that our training implementation actively clears unused variables in real time to minimize GPU memory usage. Therefore, the reported memory values correspond to the maximum observed usage within a one-minute monitoring window. All experiments were repeated three times with different random seeds (42, 64, and 128) to ensure the robustness and reproducibility of the results. The comparison results are shown in the Table 9.
Framework
Training Memory Usage
Inference Memory Usage
per-token communication
volume (Bytes)
Inference tokens
per second
FL-GLM 9810±0.00 6886±0.00 context length N × 10,335 0.75±0.03
FedSEA-LLaMA 5078±0.00 3674.00±0.00 8494 0.34±0.01
Table 9: Efficiency Comparison between FedSEA-LLaMA and FL-GLM
Appendix E Parameter Setting
FedSEA-LLaMA and Baselines
Different hyperparameters are explored to obtain the best performance of FedSEA-LLaMA and Centralized LLaMA. Specifically, we tune the LoRA rank (r) and scaling factor (alpha) as pairs, trying combinations including (4, 32), (8, 32), (16, 32), (32, 32), and (8, 16). For the learning rate (lr), values of 1e-5, 2e-5, and 3e-5 are tested. The final parameter settings were selected based on the model’s validation performance. The final parameters are shown in Table 10.
CoQA XSum MultiRC Record
FedSEA-LLaMA Lora r & alpha $(4,32)$ $(16,32)$ $(32,32)$ $(32,32)$
lr 1e-4 3e-5 1e-5 2e-5
Centralized-LLaMA Lora r & alpha $(8,16)$ $(32,32)$ $(32,32)$ $(8,32)$
lr 5e-5 3e-5 1e-5 3e-5
Table 10: Hyperparameters for FedSEA-LLaMA and Baselines
Different Noise Scales and Layers of $Block_{A}$ Experiment
We use a LoRA rank of 8 with a scaling factor of 16, a learning rate of 1e-5, and FP16 precision to investigate the effects of different noise scales and the number of blocks in $Blocks_{A}$ on downstream performance.
Ablation Study on Attention-Mask
The context length is set to 500, and the precision is set to FP16.
Ablation Study on KV Cache
The precision is set to FP16. Different context lengths are set to investigate the inference speed of FedSEA-LLaMA and FL-GLM.
Different Model Partition
All experimental results were obtained after 8,000 training steps with LoRA rank set to 8, scaling factor set to 16, a batch size of 1, FP16 precision, and a learning rate of 1e-5.

### v4

@@ Introduction
While the success of Large Language Models (LLMs) has largely stemmed from their ability to process vast amounts of public data (kaplan2020scaling), the true potential of these models lies in their application to high-quality private data. Private datasets often contain more granular and specific insights about individuals or organizations, typically collected through more controlled and precise methodologies (yang2024federated). Additionally, private data tend to be more timely, targeted, and unique, thus providing models with semantic contexts and environmental settings that more closely mirror real-world scenarios (ye2024openfedllm). Therefore, effectively incorporating private data will further advance the performance of LLMs.
@@
The distributed private data silos and the significant computational demands of LLMs pose substantial challenges for their deployment in federated environments. Private data is often isolated in local silos, such as mobile devices (ghimire2022recent), enterprise servers (huang2024federated), or medical institutions (yuan2024decentralized), as centralized storage and processing of such data may raise privacy concerns and regulatory risks (chen2024federated). Federated Learning (FL) (stremmel2021pretraining) offers a promising solution by enabling local training on user devices and only sending model parameters for aggregation. They often leverage Adapters (houlsby2019parameter) or LoRA (hu2022lora) as transmitted parameters, or employ collaborative training between large and small models (hinton2015distilling) to fine-tune LLMs. However, in extremely resource-constrained scenarios (e.g., less than 3.5GB of GPU memory), fine-tuning the LLMs remains infeasible. Therefore, how to effectively deploy LLMs in FL frameworks while ensuring privacy and scalability in resource-constrained scenarios remains an open research problem.
To address this, transformer-based federated splitting models (zheng2024safely; su2024titanic) have been proposed to adapt LLM fine-tuning to resource-constrained scenarios. Titanic (su2024titanic) handles limited server resources by partitioning LLM across clients and enabling collaborative training. FL-GLM (zheng2024safely) offloads most model parameters to a resource-ample server while retaining the input and output layer on the client. As shown in Figure 1a, FL-GLM divides Transformers into three parts, i.e., the first input block and the last output block on the client, while the middle blocks on the server. The forward propagation (FP) path is client-to-server-to-client with peer-to-peer key encryption. After the client calculates loss, gradients backpropagate (BP) from client-to-server-to-client with the same encryption. Although FL-GLM provides a potentially viable solution for clients in resource-constrained scenarios, it still faces several limitations: 1) Privacy Leakage: during forward and backward propagation, activation values and gradients from the input blocks are transmitted, which can be exploited to reconstruct original input data through reverse engineering (asnani2023reverse). 2) Communication Cost: LLMs are auto-regressive, requiring sequential token generation based on prior tokens. In federated split learning, this necessitates a full forward and backward pass with client-server communication for each token, greatly increasing communication rounds and bandwidth usage. 3) Lack of Adaptability: Fixed partition points of FL-GLM lack flexibility to downstream tasks and hardware constraints, as different devices vary in their capacity to handle workloads and parameter sizes.
@@
Traditional Federated Learning (FL) frameworks face scalability challenges with large language models (LLMs) due to high communication and computational costs (mcmahan2017communication; stremmel2021pretraining; ji2019learning). Recent federated LLM frameworks (fan2023fate; kuang2024federatedscope; ye2024openfedllm) address this by using instruction tuning, Adapters, and LoRA to reduce parameter updates, lowering communication overhead. However, transmission demands often still exceed system limits. To mitigate this, gradient compression (wu2024cg; shu2024ferret) and quantization (jianhao2024promoting) techniques have been applied. Structural pruning based on Fisher Information (liu2024fisher) further narrows updates to key layers, reducing model size without performance loss. Knowledge distillation (wu2024fedbiot; fan2024fedcollm; fan2025fedmkt) offers another solution, using smaller proxy models to transfer knowledge with minimal communication. Client heterogeneity remains a major challenge (zhang2024towards), prompting methods like LoRA-based stacking (bai2024federated; wang2024flora) to align models from diverse clients and improve fairness.
Despite advances in efficient fine-tuning and parameter reduction for federated LLMs, resource-constrained clients still face challenges with independent fine-tuning, particularly under compression overhead. Split learning (thapa2022splitfed) offers a promising alternative in such settings.
@@
Split learning is an emerging decentralized training approach designed for resource-constrained settings (abedi2024fedsl; rahman2020internet; matsubara2021neural), especially mobile or GPU-limited clients. It works by dividing a neural network into sub-networks, each processed on a different device for collaborative learning. To address the training of sequential data in LLMs, FedBERT (tian2022fedbert) introduces an innovative federated learning framework designed to pre-train language models in a split architecture, addressing the challenge of limited computational capacity on client devices. Although this design enables more efficient distribution of training workloads, it comes with drawbacks, such as increased communication overhead, fixed partition points, and susceptibility to privacy threats like embedding gradient attacks.
Recently, su2024titanic extends the scope of split learning by exploring scenarios with heterogeneous resources, particularly considering a case where the server lacks sufficient computational capacity. In their approach, the LLM is automatically divided into $k$ partitions, each assigned to one of $k$ clients, while the server merely performs aggregation. Although this strategy alleviates server-side computation, it considerably increases the overall training and inference time. Furthermore, repeated transmission of hidden states increases the risk of single points of failure and introduces potential security vulnerabilities. zheng2024safely propose FL-GLM to place the input and output blocks locally on client devices, while the remaining primary model parameters are centralized on a server with ample computational resources. And then, they employ key encryption during client-server communication to prevent reverse engineering attacks from peer clients. But the peer-to-peer key encryption is not enough because the peer clients and servers can collaborate to infer the clients’ private data.
@@
We design an extreme attack scenario to evaluate the robustness of FedSEA-LLaMA. Inspired by pasquini2021unleashing, we construct a model inversion attack in a multi-client federated setting, as illustrated in Figure 3a, where a malicious client colludes with an honest-but-curious server. The malicious client shares its private data in plaintext with the server, allowing the server to train an attack model locally. The goal of the attack model is to reconstruct the original input text from the hidden states transmitted by the client. During this process, the server does not tamper with the hidden states sent to other honest clients, thereby performing a stealthy inversion attack against their private data without awareness.
@@
During backpropagation, according to (clark2017computing), the gradient of $W_{n}$ is:
@@
Model ReCoRD COPA WSC RTE BoolQ WiC CB MultiRC CoQA Xsum F1/Acc. Acc. Acc. Acc. Acc. Acc. F1 F1a/EM F1/EM rouge-1/rouge-2 $\text{T5}_{\text{large}}$ (raffel2020exploring) 85.7/85.0 78.0 84.6 84.8 84.3 71.6 96.4 80.9/46.6 - 40.9/17.3 $\text{BART}_{\text{Large}}$ (lewis2020bart) 88.3/87.8 60.0 65.4 84.5 84.3 69.0 90.5 81.8/48.0 - 45.1/22.3 $\text{GLM}_{\text{RoBERTa}}$ (glm) 89.6/89.0 82.0 83.7 87.7 84.7 71.2 98.7 82.4/50.1 - 45.5/23.5 ChatGLM-6B (zeng2022glm) 80.2/78.7 85.0 71.2 81.6 83.4 71.0 85.7 78.2/45.6 - 37.6/12.5 FL-GLM (zheng2024safely) 79.8/78.4 85.0 71.2 80.1 81.9 69.6 85.7 79.3/46.1 62.7/49.0 37.0/11.9 LLaMA2-7B (touvron2023LLaMA) 81.3/79.5 75.0 68.3 73.3 83.8 70.9 85.7 82.3/50.1 88.6/80.9 45.9/26.1 FedSEA-LLaMA 81.3/79.4 75.0 68.4 73.3 82.4 70.8 85.7 82.7/48.0 88.6/80.7 47.6/25.0
@@
  • The SuperGLUE benchmark (wang2020superglue) evaluates eight distinct NLU tasks, from common sense reasoning to fine-grained semantic interpretation.
  • The CoQA dataset (reddy2019coqa) is designed for the task of conversational question answering(QA), featuring 8k+ dialogues across seven domains. Answers are free-form text, with nearly half of the questions requiring coreference resolution and pragmatic reasoning, simulating real-world dialogue coherence.
  • The XSum dataset (narayan2018don) evaluates abstractive summarization via extreme compression of BBC news articles into single-sentence summaries. It tests the model’s ability to abstract and condense salient information across diverse topics.
@@
For conversational QA tasks, we use Exact Match (EM) and F1 score to evaluate model performance (yatskar2019qualitative; you2022end). EM measures the percentage of predictions that exactly match the ground truth, accounting for punctuation and case, while F1 reflects the harmonic mean of precision and recall, rewarding partial overlaps.
For summarization tasks, we use standard ROUGE metrics (lin2004rouge; liu2021topic; fang2022spoken), specifically ROUGE-1 and ROUGE-2, which assess unigram and bigram overlap between generated and reference summaries, capturing content coverage and fluency.
@@
We evaluate FedSEA-LLaMA based on the LLaMA2-7B model (touvron2023LLaMA), an open-source pre-trained model developed by Meta and FL-GLM. We additionally include several strong baselines: $\text{T5}_{\text{large}}$ (raffel2020exploring), $\text{BART}_{\text{Large}}$ (lewis2020bart), $\text{GLM}_{\text{RoBERTa}}$ (glm) and ChatGLM-6B (glm2024chatglm).
@@
The server is equipped with a Gigabit Ethernet card and utilizes multiple GPUs, including two NVIDIA A6000 and two NVIDIA L40. The implementation is based on the Flower (beutel2020flower) framework, which orchestrates the transfer of intermediate data between clients and the server. Due to limited computational resources, FP32 precision and a batch size of 1 are used only for the CoQA dataset. All other experiments are conducted with FP16 precision and a batch size of 2. The random seed is consistently set to 42. Fine-tuning is performed using distributed LoRA training, where $Blocks_{A}$, $Blocks_{B}$, and $Blocks_{C}$ each maintain their own LoRA adapters. For all LoRA fine-tuning experiments in this paper, the target modules are set to q_proj, v_proj, k_proj, and o_proj.